In a market demand and supply equations are:

The demand curve is given as: P = 50 - 3Q

The supply curve is given as: P = 10 + 2Q

Assuming a perfectly competitive market:

1) What is the equilibrium price and quantity?
Answer #1

In a market demand and supply equations are:

The demand curve is given as: P = 50 - 3Q

The supply curve is given as: P = 10 + 2Q

Assuming a perfectly competitive market, equilibrium is achieved when demand and supply curve intersect each other.

If demand and supply curve intersect each other at equilibrium, then,

50 - 3Q = 10 + 2Q

or, 40 = 5Q

or, Q* = 8.

Therefore, equilibrium quantity = 8.

Equilibrium price (P*) = 50 – 3*8 = 50 – 24 = 26.

Answer – The equilibrium price and quantity are 26 and 8 respectively.